# Community Discovery Algorithm Based on Spatio-temporal Graph Embedding in Dynamic Social Networks

### v1

@@ 3.4 Time Feature Extraction
Figure 3: Structure of GRU.
@@ 3.5 Community Detection
The algorithm proposed in this paper learns node representations of dynamic networks in an unsupervised manner. To ensure consistency between nodes in the embedded and real spaces, Euclidean distance is used to measure proximity, ensuring that connected nodes are placed close to each other, while unconnected nodes are positioned far apart. This can maintain the first-order (pairwise proximity between nodes) and second-order (pairwise proximity between node neighborhood structures) proximity of nodes. Specifically, for the set $S=\left\{\left(f_{x},f_{y}\right),\forall(x,y)\in E\right\}$, in the pairwise aggregation of nodes, the following objective loss function can be used
$L=\sum_{\left(f_{x},f_{y}\right)\in S}\sum_{\left(f_{x},f_{u}\right)\in S}\left[m+\left\|f_{x}-f_{y}\right\|_{2}^{2}-\left\|f_{x}-f_{u}\right\|_{2}^{2}\right]_{+},$ (10)
where $m$ is a constant representing the size of the security domain, the operator $[\cdot]_{+}=\max(\cdot,0)$ represents the hinge loss.
In large-scale social networks, most nodes are often not connected by edges, and minimizing this loss function is computationally time-consuming. Therefore, we use the negative sampling technique (Duan et al., 2021; Yang et al., 2023) to tackle with this problem. In this technique, negative samples are collected based on the degree of the node, and is then sampled based on the noise distribution $P_{n}(v)\sim d_{v}$, where $d_{v}$ represents the degree of the node $v$. The loss function is
$L$ $=\sum_{\left(f_{x},f_{y}\right)\in S}\left(\sum_{q=1}^{Q}E_{v_{q}}\sim P_{n}(v)\right)$ (11)
×$\times\left[m+\left\|f_{x}-f_{y}\right\|_{2}^{2}-\left\|f_{x}-f_{v_{q}}\right\|_{2}^{2}\right]_{+}.$
where $Q$ is the number of negative samples. After obtaining the embedding features of the dynamic network through spatiotemporal graph embedding, the SOM algorithm is used to learn the embedding information. Competitive learning clusters the input node embeddings, resulting in the final community division of the nodes. The pseudo-code of the CDA-SGE algorithm is shown in Algorithm 1.
@@
$G=\{g_{1},g_{2},\dots,g_{t},\dots,g_{T}\}$triangleright Input graph sequence
$C=\{C_{1},C_{2},\dots,C_{T}\}$triangleright Output community sequence
@@ 4.1 Experiment Setup
All algorithms in the paper are programmed using Python, and the relevant environment is configured using Anaconda. The GCN and GRU parts of the code are built based on Pytorch. The operating system of the computer is 64 bit Windows 10 system, and the hardware configuration is Lenovo Qitian M510; AMD PRO A10-8700 3.5GHz; 8GB memory; 128GB+1TB hard drive. The data set used in this experiment is a time snapshot network, including both artificial and real network data sets. In the experiment, the parameters are taken to be $N=1000,\ s=9,\ \mu\in[0.1,0.8]$. We change the value of $\mu$ to obtain 8 dynamic artificial datasets, labeled as LFR1 to LFR8. We validate the performance of CDA-SGE using four types of dynamic networks: PS, CW, DBLP5, and Brain, ranging from small to large. The experimental evaluation uses cluster indicators, including purity, standardized mutual information, heterogeneity, and homogeneity, as performance metrics.
@@ 4.2 Comparison Algorithms
The evaluation of spatiotemporal graph embedding models is performed by comparing them with traditional static graph embedding methods (SDNE(Wang et al., 2023), DeepWalk(Perozzi et al., 2014)) and dynamic graph embedding algorithms (DynAE, DynGEM)(Goyal et al., 2020). After obtaining node embeddings, K-means clustering is used to divide the nodes into communities. In this paper, purity and normalized mutual information (NMI) are used to evaluate the quality of the clustering results. The parameter settings for the relevant dynamic graph embedding algorithm are shown in Tab. 2.
@@
To evaluate the community discovery capability of the proposed algorithm, we compared the CDA-SGE-K, where K stands for the final K-means clustering, with dynamic community discovery algorithms Dyperm and FacetNet(Lin et al., 2023), the static community discovery algorithm Louvain(Blondel et al., 2023), and the dynamic community discovery method DyCTWE based on time walk graph embedding.
@@ 4.3.1 Ablation Experiments
In complex machine learning models, which often consist of multiple modules, ablation experiments are used to assess the importance of specific components by removing certain modules and rerunning the experiments. The results are then compared with those from the original model to evaluate the impact of the removed modules. We evaluate the effectiveness of dynamic community methods based on spatiotemporal graph embedding through ablation experiments. The experiments are conducted on a dynamic artificial network (LFR2) with $\mu=0.2$. We compare the full model with GRU, GCN, and GCN-GRU. The evaluation criteria include the average NMI, completeness, and homogeneity of the dynamic network, which are used to verify the superiority of the proposed algorithm. A detailed comparison is presented in Tab 3.
@@
GCN 0.748334 0.748316 0.748352
GRU 0.695722 0.695708 0.695736
GCN-GRU 0.893428 0.892967 0.893891
CDA-SGE 0.938524 0.937010 0.940042
Table 3: Comparison of Ablation Experimental Results
Figure 4: The average evaluation results of embedding models in datasets: the average purity in LFR (a), the average NMI in LFR (b), the average purity in real dataset (c), and the average NMI in real dataset (d).
As shown in Tab. 3, the community discovery results of CDA-SGE outperform other graph embedding methods, demonstrating its effectiveness in capturing complex relationships in dynamic networks. The CDA-SGE algorithm achieves superior performance across all evaluation metrics, including NMI, completeness, and homogeneity, indicating that its integrated approach to spatiotemporal graph embedding is well-suited for dynamic community discovery tasks.
The lower performance of GRU can be attributed to its primary focus on learning the temporal changes of the nodes while neglecting the spatial structural characteristics of the network. This limitation reduces its ability to fully capture the intricate dependencies between nodes, leading to suboptimal community discovery. Additionally, the random initialization of node features without fixed semantics introduces inconsistencies in the representation of nodes, further diminishing the effectiveness of the GRU.
Furthermore, the results demonstrate that the GCN-GRU method performs significantly better than GCN alone, indicating that combining spatial topology learning with temporal sequence modeling provides a more holistic understanding of the network’s evolution. The synergy of GCN and GRU enables a better representation of evolving community structures and achieves higher performance across all metrics.
@@ 4.3.2 Comparative Analysis of Embedding Methods
In Fig. 4 we present the evaluation results of various graph embedding algorithms on dynamic artificial synthesis datasets and real datasets. We first analyze the performance of the embedding algorithms on synthetic networks. From Fig. 4 and Fig. 4, it is observed that as the complexity coefficient increases, the average evaluation results for synthetic networks show a decreasing trend. This indicates that as the complexity of the network grows, the ability of the graph embedding algorithm to represent the network structure declines. The CDA-SGE-K algorithm exhibits strong representation capabilities for artificial networks, with K-means yielding superior community partitioning accuracy compared to other embedding models. This demonstrates that the spatiotemporal graph embedding model effectively captures spatiotemporal relationships and preserves node interaction information in complex dynamic networks. The dynamic graph embedding methods DynGEM and DynAE can also provide good embedding features, and the evaluation results of community partitioning achieve good results when the network complexity is low. However, as complexity increases, the evaluation results rapidly decline. This may be attributed to the fact that these two algorithms primarily focus on learning the spatial structure of nodes during graph embedding. For instance, DynGEM only considers the influence of the previous time snapshot on the current one, neglecting the long-range dependencies of dynamic nodes. Consequently, as network complexity increases, the graph embedding performance declines, impacting community discovery effectiveness. The static embedding algorithm DeepWalk achieves strong results, while SDNE shows stable but modest performance on artificial datasets, with minimal changes in NMI values. This suggests that SDNE can consistently characterize dynamic artificial networks, but its ability to extract node features requires further improvement.
@@
We then analyze the performance of the embedded model in real networks and present the results in Fig. 5. It is evident that all graph embedding models perform better in node partitioning accuracy on smaller dynamic datasets than on larger ones. The decrease in accuracy as the network size grows can be attributed to the increase in the number of nodes and the more complex interaction relationships between them. Larger networks introduce additional challenges, such as denser connections, overlapping communities, and a greater diversity of dynamic interactions, which make it harder for models to effectively capture and preserve all relevant spatiotemporal features.
@@
Despite these challenges, the CDA-SGE proposed in this paper consistently maintains high accuracy across both small and large datasets. This highlights its robustness and its ability to effectively extract dynamic spatiotemporal information, allowing it to characterize the evolution of real-world networks more accurately. The model’s ability to integrate spatial and temporal embeddings enables it to uncover subtle patterns and interactions that are often missed by other methods.
@@
Among the baseline models, DeepWalk achieves the second-best performance due to its ability to learn meaningful node representations through random walks. However, its lack of temporal modeling limits its performance on dynamic datasets. SDNE, on the other hand, yields relatively lower accuracy on large-scale dynamic datasets, probably because its emphasis on preserving local structures struggles to scale effectively with the increasing complexity of larger networks. Among dynamic graph embedding algorithms, DynGEM outperforms DynAE in community discovery, highlighting that SDNE’s embedding performance on the network surpasses that of fully connected neural networks.

### v2

@@ 3.4 Time Feature Extraction
Figure 3: Structure of the GRU used for temporal feature extraction in GTENN.
@@ 3.5 Community Detection
The algorithm proposed in this paper learns node representations of dynamic networks in an unsupervised manner. To ensure that embeddings reflect the underlying graph topology, we adopt a distance-preserving objective based on pairwise proximity: connected nodes should be embedded close to each other, while unconnected nodes should be placed farther apart. This promotes both first-order (direct link) and second-order (neighborhood structure) proximity preservation.
Let $\mathcal{S}=\{(f_{x},f_{y})\mid(v_{x},v_{y})\in E\}$ be the set of embedding pairs for connected nodes. The pairwise ranking-based loss function is defined as:
$L=\sum_{(f_{x},f_{y})\in\mathcal{S}}\sum_{(f_{x},f_{u})\in\mathcal{S}}\left[m+\|f_{x}-f_{y}\|_{2}^{2}-\|f_{x}-f_{u}\|_{2}^{2}\right]_{+},$ (12)
where $\cdot]_{+}=\max(\cdot,0)$ denotes the hinge loss function, and $m>0$ is a predefined margin constant that reflects the security domain—that is, the minimum separation distance we wish to enforce between a positive (connected) and negative (unconnected) sample. A larger value of $m$ encourages embeddings of unrelated nodes to be pushed further apart.
In large-scale graphs, evaluating all unconnected node pairs is computationally prohibitive. To alleviate this, we adopt the negative sampling[15, 17] strategy. Instead of exhaustively computing losses over all non-edges, we sample a small number of unconnected node pairs as negative examples for each positive pair $(v_{x},v_{y})$. Specifically, negative samples $v_{q}$ are drawn from a noise distribution $P_{n}(v)$, which we define proportionally to node degree: $P_{n}(v)\sim d_{v}^{3/4}$. The modified loss function becomes:
$L=\sum_{(f_{x},f_{y})\in\mathcal{S}}\left(\sum_{q=1}^{Q}\mathbb{E}_{v_{q}\sim P_{n}(v)}\left[m+\|f_{x}-f_{y}\|_{2}^{2}-\|f_{x}-f_{v_{q}}\|_{2}^{2}\right]_{+}\right),$ (13)
where $Q$ is the number of negative samples drawn per positive pair.
After learning the spatial-temporal embeddings $f_{t}^{i}$ of nodes in each snapshot, we apply the Self-Organizing Map (SOM) to discover latent community structure. SOM is an unsupervised neural clustering algorithm that maps high-dimensional vectors into a low-dimensional (typically 2D) grid while preserving topological relations. Using competitive learning, SOM identifies clusters of nodes with similar structural embeddings, forming the final community partitions. The full procedure is summarized in Algorithm 1.
@@
$G=\{g_{1},g_{2},\dots,g_{t},\dots,g_{T}\}$ $\triangleright$ Input graph sequence
$C=\{C_{1},C_{2},\dots,C_{T}\}$ $\triangleright$ Output community sequence
@@ 4.1 Experiment Setup
All algorithms in this study are implemented in Python 3.10, with the GCN and GRU components developed using PyTorch 2.4.1. Experiments are conducted on a Hygon C86 32-core CPU machine with a Kylin 10 Linux-based operating system, equipped with 512 GB of memory and 8 Hygon DCU GPUs.
The datasets used in this study consist of time-snapshot dynamic networks, including both synthetic and real-world network datasets. For the synthetic data, we adopt the widely used LFR (Lancichinetti–Fortunato–Radicchi) benchmark model[19], which generates graphs with built-in community structure. In this simulation: $N=1000$ specifies the total number of nodes in the graph; $s=9$ defines the number of time snapshots (i.e., network evolution stages); and $\mu\in[0.1,0.8]$ is the mixing parameter, which controls the community strength: lower values of $\mu$ indicate clearer communities (most edges remain within communities), while higher values increase the inter-community connections, making communities harder to detect. By varying $\mu$, we generate eight synthetic dynamic network datasets labeled from LFR1 to LFR8. These provide different levels of community detectability over time.
To evaluate our GTENN model, we also apply it to four real-world dynamic networks of varying size and complexity: PS network representing face-to-face interactions in educational settings [20], CSMN network capturing online social interactions in college environments [21], DBLP collaboration network from computer science publications [20], and Brain networks derived from neuroimaging data [22]. Cluster indicators were used as evaluation criteria in the experiment, mainly including purity, standardized mutual information, heterogeneity, and homogeneity.
The main parameters that need to be estimated in this algorithm include the number of GCN layers Land the weight matrices $W^{l}_{t}$ for each layer, the GRU weight matrices $U$ and $V$ along with bias matrices $B$, the embedding dimension $d$, the number of competing units in the SOM network along with neighborhood range and learning rate, and the number of negative samples $Q$ and security domain size $m$. These parameters are learned through the training process, where the neural network parameters are optimized via backpropagation while the SOM parameters are configured based on competitive learning principles.
4.2 Evaluation Metrics
The experimental evaluation employs four clustering quality metrics to assess community discovery performance:
Purity: Measures the proportion of correctly classified nodes within each cluster, calculated as
$\text{Purity}=\frac{1}{N}\sum_{k=1}^{K}\max_{j}|C_{k}\cap T_{j}|$ (14)
where $C_{k}$ represents the $k$-th discovered community and $T_{j}$ represents the $j$-th true community. Higher purity values indicate better clustering quality.
Normalized Mutual Information (NMI): Quantifies the agreement between discovered and ground-truth communities, defined as
$\text{NMI}=\frac{2\cdot I(C,T)}{H(C)+H(T)}$ (15)
where $I(C,T)$ is the mutual information and $H(\cdot)$ denotes entropy. NMI ranges from 0 to 1, with values closer to 1 indicating better performance.
Homogeneity: Measures whether each cluster contains only members of a single class, ensuring internal consistency within discovered communities.
Completeness: Evaluates whether all members of a given class are assigned to the same cluster, assessing the ability to group related nodes together.
4.3 Datasets
In this experiment, synthetic and real networks are used to comprehensively evaluate algorithm performance. The LFR benchmark networks provide controllable community structures for systematic testing. Real datasets include: PS, CSMN, Brain and DBLP. The summary of these datasets structures can be found in 2
4.3.1 Synthetic Datasets
We used the LFR benchmark to generate dynamic artificial datasets. The LFR benchmark is widely recognized as a standard for evaluating community detection algorithms because it can generate networks with controllable community structure and realistic properties. The following parameters were used:
  • $N=1000$: Total number of nodes in the network
  • $s=9$: Number of time snapshots in the dynamic sequence
  • $\mu\in[0.1,0.8]$: Mixing parameter that controls network complexity
  • Average degree: $\langle k\rangle=15$
  • Maximum degree: $k_{max}=30$
  • Community size range: $[10,50]$ nodes per community
  • Power-law exponent for degree distribution: $\gamma=2.5$
  • Power-law exponent for community size distribution: $\beta=1.5$
The mixing parameter $\mu$ determines the fraction of edges that connect nodes to communities other than their own, where smaller $\mu$ values indicate clearer community structure and larger $\mu$ values represent more ambiguous community boundaries. By varying only the value of $\mu$, we obtained 8 dynamic artificial datasets denoted as LFR1 ($\mu=0.1$), LFR2 ($\mu=0.2$), LFR3 ($\mu=0.3$), LFR4 ($\mu=0.4$), LFR5 ($\mu=0.5$), LFR6 ($\mu=0.6$), LFR7 ($\mu=0.7$), and LFR8 ($\mu=0.8$), corresponding to different levels of community structure clarity from very clear to highly ambiguous.
4.3.2 Real-world Datasets
To validate the performance of GTENN on real-world scenarios, we employed four types of dynamic networks with varying scales and characteristics:
PS (Primary School) Dataset: This dataset records the face-to-face interactions between students and teachers in a primary school in Lyon, France, collected over two days using wearable sensors. The network contains 242 nodes (students and teachers) and evolves over 125 time steps with 20-second intervals. The dataset captures the natural formation and dissolution of social groups in an educational environment, making it ideal for studying temporal community dynamics.
CSMN (College Social Message Network) Dataset: This dataset represents online social interactions among college students over a semester period. It contains 1,899 nodes representing students and 59,835 temporal edges representing various types of interactions including messages, wall posts, and friendship formations. The network evolves over 193 time steps, each representing one day. This dataset is characterized by periodic patterns corresponding to academic schedules and social events.
DBLP Dataset: This is a collaboration network derived from the DBLP computer science bibliography database, focusing on publications from 2000 to 2005. The network contains 12,334 authors as nodes and co-authorship relationships as edges. The temporal evolution spans 6 years, with each time step representing one year. The dataset exhibits clear research community structures that evolve as researchers change their collaboration patterns and research interests.
Brain Dataset: This dataset represents functional brain networks derived from fMRI data, where nodes correspond to brain regions and edges represent functional connectivity between regions. The network contains 264 nodes (brain regions) and evolves over 200 time steps, each representing a different time window during a cognitive task. This dataset is particularly challenging due to its high-dimensional nature and complex temporal dynamics.
Dataset Nodes Time Steps Avg. Edges/Step Domain
LFR1-8 1,000 9 $\sim$7,500 Synthetic
PS 242 125 $\sim$400 Social
CSMN 1,899 193 $\sim$310 Social
DBLP 12,334 6 $\sim$25,000 Collaboration
Brain 264 200 $\sim$1,200 Biological
Table 2: Statistics of the dynamic network datasets used in our experiments. Each dataset is characterized by the number of nodes, time steps, average edges per step, and its corresponding domain.
@@ 4.4 Comparison Algorithms
In order to evaluate the ability of the algorithm for community discovery in this article, it is mainly compared with dynamic community discovery algorithms Dyperm and FacetNet, static community discovery algorithm Louvain, dynamic community discovery method DyCTWE based on time walk graph embedding, and GTENN-K using K-means as the final clustering method. This algorithm is referred to as GTENN in this paper, and purity and NMI are used to evaluate the quality of clustering results.
The evaluation of spatiotemporal graph embedding models is performed by comparing them with traditional static graph embedding methods SDNE[23], DeepWalk[24] and dynamic graph embedding algorithms DynAE, DynGEM[25]. After obtaining node embeddings, K-means clustering is used to divide the nodes into communities. In this paper, purity and normalized mutual information (NMI) are used to evaluate the quality of the clustering results. The parameter settings for the relevant dynamic graph embedding algorithm are shown in Tab. 3.
@@
To evaluate the community discovery capability of the proposed algorithm, we compared the GTENN-K, where K stands for the final K-means clustering, with dynamic community discovery algorithms Dyperm and FacetNet[26], the static community discovery algorithm Louvain[27], and the dynamic community discovery method DyCTWE based on time walk graph embedding.
@@ 4.5.1 Ablation Experiments
In complex machine learning models, which often consist of multiple modules, ablation experiments are used to assess the importance of specific components by removing certain modules and rerunning the experiments. The results are then compared with those from the original model to evaluate the impact of the removed modules. We evaluate the effectiveness of dynamic community methods based on spatiotemporal graph embedding through ablation experiment. This experiment is conducted on a dynamic artificial network (LFR2) with $\mu=0.2$. We compare the full model with GRU, GCNs, and GCNs-GRU. The evaluation criteria include the average NMI, completeness, and homogeneity of the dynamic network, which are used to verify the superiority of the proposed algorithm. A detailed comparison is presented in Tab 4.
As shown in Tab. 4, the proposed GTENN consistently outperforms its individual components, demonstrating its effectiveness in capturing the complex spatiotemporal relationships of dynamic networks. GTENN achieves superior performance across all evaluation metrics, including NMI, completeness, and homogeneity, confirming that its integrated design is well-suited for dynamic community discovery. The relatively low performance of the GRU baseline can be attributed to its emphasis on temporal dynamics while neglecting spatial structural characteristics. Without explicit modeling of network topology, GRU struggles to capture node dependencies, leading to weaker community detection. In addition, the random initialization of node features without fixed semantics introduces inconsistencies in representation, further diminishing its effectiveness. Furthermore, the GCNs-GRU hybrid achieves substantially better results than GCNs alone, highlighting the importance of jointly modeling spatial topology and temporal evolution. This synergy between GCNs and GRU provides a more comprehensive representation of evolving communities and yields consistently higher performance across all metrics.
@@
GCNs 0.748226 0.748214 0.748249
GRU 0.705708 705697 0.705726
GCNs-GRU 0.902130 0.900143 0.902308
GTENN 0.948902 0.948017 0.949921
Table 4: Comparison of ablation experimental results on LFR2 dataset ($\mu=0.2$). The table demonstrates the effectiveness of different component combinations in GTENN
@@ 4.5.2 Comparative Analysis of Embedding Methods
In Fig. 4 we present the evaluation results of various graph embedding algorithms on dynamic artificial synthetic datasets and real datasets. We first analyze the performance of the embedding algorithms on synthetic networks.
(a) Average purity on Synthetic Datasets
(b) Average NMI on Synthetic Datasets
(c) Average Purity on Real Datasets
(d) Average NMI on Real Datasets
Figure 4: Performance comparison of graph embedding methods on synthetic and real datasets. Results show that GTENN consistently achieves higher purity and NMI scores compared to static methods (SDNE, DeepWalk) and dynamic methods (DynAE, DynGEM).
@@
From Fig. 4(a) and  4(b), it is observed that as the complexity coefficient increases, the average evaluation results for synthetic networks show a decreasing trend. This indicates that as the complexity of the network grows, the ability of the graph embedding algorithm to represent the network structure declines. The GTENN-K algorithm exhibits strong representation capabilities for artificial networks, with K-means yielding superior community partitioning accuracy measured in purity and NMI compared to other embedding models. This demonstrates that the spatiotemporal graph embedding model effectively captures spatiotemporal relationships and preserves node interaction information in complex dynamic networks. The dynamic graph embedding methods DynGEM and DynAE can also provide good embedding features, and the evaluation results of community partitioning achieve good results when the network complexity is low. However, as complexity increases, the evaluation results rapidly decline. This may be attributed to the fact that these two algorithms primarily focus on learning the spatial structure of nodes during graph embedding. For instance, DynGEM only considers the influence of the previous time snapshot on the current one, neglecting the long-range dependencies of dynamic nodes. Consequently, as network complexity increases, the graph embedding performance declines, impacting community discovery effectiveness. The static embedding algorithm DeepWalk achieves strong results, while SDNE shows stable but modest performance on artificial datasets, with minimal changes in NMI values. This suggests that SDNE can consistently characterize dynamic artificial networks, but its ability to extract node features requires further improvement.
We then analyze the performance of the embedded model in real networks and present the results in Fig. 4(c) and  4(d).
@@
From Fig. 4(c) and  4(d), it is evident that all graph embedding models perform better in node partitioning accuracy measured in purity and NMI on smaller dynamic datasets than on larger ones. The decrease in purity and NMI as the network size grows can be attributed to the increase in the number of nodes and the more complex interaction relationships between them. Larger networks introduce additional challenges, such as denser connections, overlapping communities, and a greater diversity of dynamic interactions, which make it harder for models to effectively capture and preserve all relevant spatiotemporal features.
@@
Despite these challenges, the GTENN proposed in this paper consistently maintains high accuracy measured in purity and NMI across both small and large datasets. This highlights its robustness and its ability to effectively extract dynamic spatiotemporal information, allowing it to characterize the evolution of real-world networks more accurately. The model’s ability to integrate spatial and temporal embeddings enables it to uncover subtle patterns and interactions that are often missed by other methods.
@@
Among the baseline models, DeepWalk achieves the second-best performance due to its ability to learn meaningful node representations through random walks. However, its lack of temporal modeling limits its performance on dynamic datasets. SDNE, on the other hand, yields relatively lower purity and NMI on large-scale dynamic datasets, probably because its emphasis on preserving local structures struggles to scale effectively with the increasing complexity of larger networks. Among dynamic graph embedding algorithms, DynGEM outperforms DynAE in community discovery, highlighting that SDNE’s embedding performance on the network surpasses that of fully connected neural networks.